HUBBLE SPACE TELESCOPE DAILY REPORT #1997 PERIOD COVERED: 0000Z (UTC) 11/03/97 - 0000Z (UTC) 11/04/97 Daily Status Report as of 308/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.3 Completed Four Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of NIC/1/2 7454 (The Stellar Content of Powerful Radio Galaxies -- A NICMOS Survey of 0.8 < z < 1.8 3CR Sources) The Near Infrared Camera (Cameras One and Two) was used to make survey observations of the powerful radio galaxies 3C184 and 3C368. The study of distant radio galaxies is complicated by their powerful active nucleus, whose emission dominates at rest-frame ultraviolet wavelengths, producing the observed peculiar, elongated morphologies; the underlying stellar component is lost in the glare. The observations were completed as planned, and no anomalies were reported. 1.5 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Four Sets of STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.7 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy 1148+56W1 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CC 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/1/2 7321 (Spatially Resolved Multispectral Investigation of Titan's Troposphere and Stratosphere) The Near Infrared Camera (Cameras One and Two) was used to make observations of Saturn's moon Titan in order to study its atmosphere. Saturn's moon Titan has a thick, hazy atmosphere obscuring a surface that has only recently been observed in the near-infrared. Despite observations of surface albedo features, the nature of the surface remains hidden due to our imprecise knowledge of the atmosphere. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2/NIC/1/2/3 7414 (Imaging of a New Beta Pic-like Circumstellar Dust Disk) The WF/PC-2 and the Near Infrared Camera (Cameras One, Two, and Three) were used to obtain high-resolution mapping of the recently discovered circumstellar disk around the B5V binary BD+31D643 (a.k.a., HD281159). The observations completed as scheduled, with no reported anomalies. 1.11 Completed Four Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.12 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 6467 (Precise Astrometry in the Core of the Globular Cluster 47 Tuc: A Complete Census of High-Velocity Stars) The WF/PC-2 was used to make images of the cluster NGC-104 (a.k.a., 47 Tuc) in order to make precise astrometric observations of the clusters core. Hard binaries are known to exist in cluster cores, e.g., in the form of millisecond pulsars. The observations were completed as planned, and no anomalies were reported. 1.14 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell V4332-SGR in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 As described in HSTAR 6231, one acquisition defaulted to fine lock on one FGS only at 308/0811Z. This may have affected one STIS and two NICMOS proposals which will be identified in tomorrow's report. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: The STIS error counter was reset at 307/1523Z. The STIS bad pixel table was uplinked at 307/2213Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH